# Getting Explicit Instruction Right

### v1

@@ 5 Results
For each of our four outcomes, we construct a distinct synthetic control, with its associated set of weights. [^2] We produce our estimates and plots using the scpi R package (Cattaneo et al., 2022). However, since the four sets of weights estimated are identical to each other after rounding them to two decimals, we treat them as one in our discussion. As shown in Table 1, Synthetic Charlestown South’s biggest “donor schools” are the Charlestown (38%) and Hillsborough (36%) public schools—both in the vicinity of Charlestown South. This is because one of the features we matched on was “radial proximity to Charlestown South”.
@@
The results of our analysis are summarized in Figures 1 and 2. The patterns are similar across topics and are stable across different estimators, as shown in Figures 4, 5, 6 and 7 in the Appendix.
@@
In light of this, we focus on the effect strictly after three years from the introduction of Explicit Instruction and find an average effect of 75.66 points (or 2.72 standard deviations) on Year-3 Numeracy scores and of 87.50 NAPLAN points (or 2.55 s.d) on Reading scores (see Figures 1(a) and 2(a)). Even if we want to calculate the treatment effect on Year-3 NAPLAN scores starting from the year when Explicit Instruction was first introduced, its impact remains as large as 44.66 points (or 1.60 standard deviations) in Numeracy and 49.22 points (or 1.44 s.d.) in Reading.
The patterns in NAPLAN scores are also similar for Year-5 students across topics (see Figures 1(b) and 2(b)). Here, the effect of Explicit Instruction seems to start sooner—one or two years after the approach is introduced—but the magnitude is substantially smaller. If we continue to focus on the average treatment effect strictly three years after the treatment date, we find effects of less than half in size compared to Year-3 impacts—31.90 points (or 1.14 s.d.) in Numeracy and 40.38 points (or 1.31 s.d.) in Reading. As above, even if we start calculating the effects for Year-5 students immediately after the pedagogy was officially introduced, we find large effects of 26.10 points (or 0.93 s.d.) in Numeracy and 36 points (or 1.17 s.d.) in Reading.
(a) Year-3
(b) Year-5
Figure 1: Effect of Explicit Instruction on NAPLAN Numeracy scores
(a) Year-3
(b) Year-5
Figure 2: Effect of Explicit Instruction on NAPLAN Reading scores
@@
To further test the sensitivity of our estimates, we take a leave-one-out approach. Specifically, we select the top 8 donor schools by weight and run our main specification 8 times, each time excluding one of these schools from the donor pool. As shown in Table 4 and Figures 8 and 9 in the appendix, the results are remarkably similar to our main estimates—both when taking each of the 8 sets of estimates individually (Figures 8 and 9) and when looking at their aggregate average treatment effect (Table 4). This is evidence that no individual donor school is driving the results.
@@
Finally, our results are neither sensitive to the exclusion of covariates nor to the choice of synthetic control estimator. We show this in the Appendix, Figures 4, 5, 6 and 7. In panel (a) of each figure, we plot the results using the Abadie et al. (2010) estimator without using any covariates. In the (b) panels, we use the Ferman and Pinto (2021) estimator, which is a modification of Abadie et al. (2010) where the pretreatment average is calculated for all units and the data demeaned. In the (c) panels, we use the Hsiao et al. (2012) estimator, where weight vectors $\mathbf{W}$ is estimated via ordinary least squares, and are hence not constrained to sum up to one. In the (d) panels, we use the Chernozhukov et al. (2021) estimator, where weight vector $\mathbf{W}$ is estimated via constrained Lasso, hence relaxing the constraint on $\mathbf{W}$ compared to Abadie et al. (2010) [^4] The Chernozhukov et al. (2021) estimator allows $\mathbf{W}$ to be any vector with bounded $\ell_{1}$-norm.. Our results remain remarkably consistent.
6 Conclusion
@@
We analyze the effect of correctly implementing Explicit Instruction on learning outcomes, as measured by standardized test scores. We focus on the case of primary public school Charlestown South, in New South Wales. We find very large effects in both Numeracy and Reading, particularly on Year-3 students. Our sensitivity analysis provides evidence that the results are not an artifact either of our preferred method or of our preferred model specification.

### v2

@@ 4 Results
For each of our four outcomes, we construct a distinct synthetic control, with its associated set of weights. [^4] We produce our estimates and plots using the scpi R package (41). However, since the four sets of weights estimated are identical to each other after rounding them to two decimals, we treat them as one in our discussion. As shown in Table 1, Synthetic Charlestown South’s biggest “donor schools” are the Charlestown (38%) and Hillsborough (36%) public schools—both in the vicinity of Charlestown South. This is because one of the features we matched on was “radial proximity to Charlestown South”.
@@
The results of our analysis are summarized in Figure 1. The patterns are similar across topics and are stable across different estimators, as shown in Figures 3, 4, 5 and 6 in the Appendix.
@@
In light of this, we focus on the effect strictly after three years from the introduction of Explicit Instruction and find an average effect of 75.66 points (or 2.72 standard deviations) on Year-3 Numeracy scores and of 87.50 NAPLAN points (or 2.55 s.d) on Reading scores (see Figures 1(a) and 1(c)). Even if we want to calculate the treatment effect on Year-3 NAPLAN scores starting from the year when Explicit Instruction was first introduced, its impact remains as large as 44.66 points (or 1.60 standard deviations) in Numeracy and 49.22 points (or 1.44 s.d.) in Reading.
The patterns in NAPLAN scores are also similar for Year-5 students across topics (see Figures 1(b) and 1(d)). Here, the effect of Explicit Instruction seems to start sooner—one or two years after the approach is introduced—but the magnitude is substantially smaller. If we continue to focus on the average treatment effect strictly three years after the treatment date, we find effects of less than half in size compared to Year-3 impacts—31.90 points (or 1.14 s.d.) in Numeracy and 40.38 points (or 1.31 s.d.) in Reading. As above, even if we start calculating the effects for Year-5 students immediately after the pedagogy was officially introduced, we find large effects of 26.10 points (or 0.93 s.d.) in Numeracy and 36 points (or 1.17 s.d.) in Reading.
(a) Numeracy Year-3
(b) Numeracy Year-5
(c) Reading Year-3
(d) Reading Year-5
Figure 1: Effect of Explicit Instruction on NAPLAN scores
@@
To further test the sensitivity of our estimates, we take a leave-one-out approach. Specifically, we select the top 8 donor schools by weight and run our main specification 8 times, each time excluding one of these schools from the donor pool. As shown in Table 4 and Figures 7 and 8 in the appendix, the results are remarkably similar to our main estimates—both when taking each of the 8 sets of estimates individually (Figures 7 and 8) and when looking at their aggregate average treatment effect (Table 4). This is evidence that no individual donor school is driving the results.
@@
Finally, our results are neither sensitive to the exclusion of covariates nor to the choice of synthetic control estimator. We show this in the Appendix, Figures 3, 4, 5 and 6. In panel (a) of each figure, we plot the results using the Abadie et al. (42) estimator without using any covariates. In the (b) panels, we use the Ferman and Pinto (43) estimator, which is a modification of Abadie et al. (42) where the pretreatment average is calculated for all units and the data demeaned. In the (c) panels, we use the Hsiao et al. (44) estimator, where weight vectors $\mathbf{W}$ is estimated via ordinary least squares, and are hence not constrained to sum up to one (see appendix for more methodological context). In the (d) panels, we use the Chernozhukov et al. (45) estimator, where weight vector $\mathbf{W}$ is estimated via constrained Lasso, hence relaxing the constraint on $\mathbf{W}$ compared to Abadie et al. (42) [^6] The Chernozhukov et al. (45) estimator allows $\mathbf{W}$ to be any vector with bounded $\ell_{1}$-norm.. Our results remain remarkably consistent.
5 Discussion
As mentioned, our estimated treatment effects range between 0.9 and 2.7 s.d. in size and apply to a median-performing school. According to Kraft (46)’s review of 747 RCTs, this places our intervention in the 99th percentile of education interventions by effect size. For context, a median intervention has an impact of 0.10 s.d. and a 90th percentile intervention of 0.40 s.d. (46).
Given the impact size alone of our intervention, there might be concerns that it is the leadership of the principal, rather than Explicit Instruction itself driving this radical change. This is only correct in the sense that it was the principal who identified that peer-modelling a top explicit-instruction school could turn his school into a high-performing one. I present evidence supporting this view below.
First, the principal started working at Charlestown South in 2012, while the change to Explicit Instruction started in 2014. Leadership and other principal-specific effects would have likely materialised before 2014 if they were meaningful. Second—when interviewed for another research project—the principal did not report any other intervention being introduced during his tenure. Third, word of his approach of peer-modelling a successful explicit-instruction school spread among other schools, which started sending staff to model Charlestown South. Indeed, the demand for learning Charlestown South approach directly became such that in 2019 the principal set up a termly “EI Open Day” dedicated to showing to the principals and staff from other schools the Explicit Instruction approach in action (47; 48; and see Facebook post 49, for an example, also reported in Figure 9 in the appendix).
Not only the teaching community believes this specific implementation of Explicit Instruction is ideal and replicable—state governments believe it too. Indeed, in 2024, the Victorian state government tasked at LaTrobe University (50) to link successful Explicit Instruction schools with schools that have not adopted this pedagogy yet. This will support the state-wide transition to structured phonics and explicit instruction among students from the first year of compulsory schooling (Prep) to grade two.
Finally, if we had to identify one key channel through which Explicit Instruction had an impact, it would be literacy, math and writing warm-ups. Indeed, both the high-performing Melbourne school and—after switching to Explicit Instruction—Charlestown South credit the warm-ups as the main driver of student performance, thanks to their ability to move information acquired in the previous lesson from the short-term to the long-term memory. The “warm-up” (or “Daily Review”) is a 5-20-minute routine scheduled at the start of a lesson that rehearses past literacy, mathematics and writing content, and checks prerequisite skills, hence preparing students to learn new material (51, 52). While the exact design of the warm-up depends on the grade of the students, during literacy warm-ups students might engage in quick exercises like revisiting letter sounds, practicing common words, or constructing simple sentences. In writing warm-ups, teachers may review sentence structures, practice punctuation, or brainstorm vocabulary. These activities are aimed at strengthening core reading and writing skills, and prepare students for more complex writing tasks. In mathematics warm-ups, teachers may prompt students with rapid questions on number facts, counting sequences, or basic calculations. This helps solidify numerical understanding and mental math skills. These warm-ups are interactive, often involving the whole class, making learning more effective and engaging (see 53, 54, for detailed recent materials on this topic).
Concluding remarks.
@@ 5 Discussion
We analyze the effect of correctly implementing Explicit Instruction on learning outcomes, as measured by standardized test scores. We focus on the case of primary public school Charlestown South, in New South Wales. We find large effects in both Numeracy and Reading, particularly on Year-3 students. Our sensitivity analysis provides evidence that the results are not an artifact either of our preferred method or of our preferred model specification.
@@
6 Materials and Methods
To create a synthetic Charlestown South, first, we need to select a pool of schools that we judge as being similar to Charletown South, but were not subject to the same intervention—here, learning the Explicit Instruction pedagogy by peer-modelling a successful Explicit Instruction school. Among all primary public NSW schools, we selected schools that have some key features that are identical to Charlestown South. We exact-match these features as we see them as too important to be matched using the minimum-distance approach embedded in SCM (see the next paragraph). They have the same level of remoteness (55), “Major Cities of Australia”; they are co-educational; they cover Kindergarten to Year 6, and their student body comprises between 10-15% of students with a language background other than English (Charlestown South’s is 12%). We are left with 108 schools which, in the language of synthetic control, are referred to as donor schools.
Fourth, we construct a synthetic control. This involves creating a “synthetic” version of Charlestown South as similar as possible as the real Charlestown South. Since we cannot observe Charlestown South had it not been treated in 2014, we use its synthetic version to approximate this counterfactual scenario. This allows us to estimate the treatment effect of implementing Explicit Instruction via peer-modelling over time—by comparing Charlestown South after treatment with its synthetic version. The synthetic control is built by finding a “mix” (the convex hull) of donor schools such that associated donor school weights (i) leading to pre-treatment paths of NAPLAN scores “most similar” to Charlestown South’s and (ii) produce school features of synthetic Charlestown South that are “most similar” to those of the real one.